(b) what is the net charge of all these electrom? nser han a man per mde d 1.0gmd -noh-at mule de0%O) has 10 eledre ns. (a) iso man .lectrons are there win in of walir? (i liter is endvalent to 1.00 ×10,㎡ (a) Number (b) Number units (No un Units Question Attempts: 0 ed 3 uned Can someone help me answer this question
Answer #1

Solution,

No. of electrons = 10 x 6.022 x 10^23 x 1 = 6.022 x 10^24 electrons

A)

Mass, m = 1000 x 3.89/1000 = 3.89 kg = 3890 g

No. of electrons = (6.022 x 10^24 x 3890)/18 = 1.30 x 10^27 electrons

B)

Required charge, Q = 1.30 x 10^27 x - 1.6 x 10^-19 = -2.08 x 10^8 C

Comment in case any doubt please rate my answer....